### v4

# OPeRA: A Dataset of Observation, Persona, Rationale, and Action for Evaluating LLMs on Human Online Shopping Behavior Simulation

Abstract
Can large language models (LLMs) accurately simulate the next web action of a specific user? While LLMs have shown promising capabilities in generating “believable” human behaviors, evaluating their ability to mimic real user behaviors remains an open challenge, largely due to the lack of high-quality, publicly available datasets that capture both the observable actions and the internal reasoning of an actual human user. To address this gap, we introduce OPeRA, a novel dataset of Observation, Persona, Rationale, and Action collected from real human participants during online shopping sessions. OPeRA is the first public dataset that comprehensively captures: user personas, browser observations, fine-grained web actions, and self-reported just-in-time rationales. We developed both an online questionnaire and a custom browser plugin to gather this dataset with high fidelity. Using OPeRA, we establish the first benchmark to evaluate how well current LLMs can predict a specific user’s next action and rationale with a given persona and <observation, action, rationale> history. This dataset lays the groundwork for future research into LLM agents that aim to act as personalized digital twins for human  [^1] https://huggingface.co/datasets/NEU-HAI/OPeRA.
[…]
booktabs
OPeRA: A Dataset of Observation, Persona, Rationale, and Action for Evaluating LLMs on Human Online Shopping Behavior Simulation
Ziyi Wang1, Yuxuan Lu1, Wenbo Li2, Amirali Amini1, Bo Sun1, Yakov Bart1, Weimin Lyu3, Jiri Gesi 4, Tian Wang4 Jing Huang4, Yu Su5, Upol Ehsan1, Malihe Alikhani1, Toby Jia-Jun Li6, Lydia Chilton7, Dakuo Wang1, 1Northeastern University, 2University of Southern California, 3Stony Brook University, 4Independent Researcher, 5Ohio State University, 6University of Notre Dame, 7Columbia University Correspondence: wang.ziyi19@northeastern.edu, d.wang@northeastern.edu
1 Introduction
Figure 1: We developed ShoppingFlow plugin (Figure 2) to collect user shopping behavior over a four-week period, resulting in OPeRA-full dataset. This dataset comprises 692 sessions from 51 unique users, containing 28,904 real-user <action, observation> pairs and 604 user-annotated rationales (Figure 3). After postprocessing, we obtained OPeRA-filtered, which includes 527 sessions, 5,856 <action, observation> pairs, and 207 rationales. We then benchmarked four LLMs’ performance on user next action prediction task, results in Table 6).
Large language model (LLM) agents have exhibited impressive performance across diverse tasks, including planning, reasoning, and acting in web-based environments Xie et al. (2024); Yao et al. (2023); Jin et al. (2025). A promising frontier in this area is human behavior simulation, where LLM agents generate user-like action sequences on digital platforms Chen et al. (2025). These agents (i.e., role-playing agents) are increasingly used in applications such as UI/UX testing Lu et al. (2025b), social science research Park et al. (2024), accessibility testing Taeb et al. (2024), and personal digital assistants OpenAI (2025b). Yet, while these agents can generate believable behavior, the more critical question remains: can they generate behavior that accurately aligns with real human?
Despite progress in agent-based behavior simulation Park et al. (2023, 2024), current works still have limitations. First, most existing evaluations focus on aggregate outcomes (e.g., survey responses or end-task completions). These methods overlook the step-wise rationale and actions that underlie user behavior patterns Chen et al. (2025). For example, Park et al. (2024) compared LLM agents’ survey results with real humans by replicating various social science studies. Furthermore, many role-play agents rely solely on prompting without grounding in real human data training, which limits their accuracy and personalization. Although some recent efforts incorporate user behavior data via fine-tuning Lu et al. (2025a), these datasets are often proprietary or lack critical detail, such as the reasoning or persona behind user actions.
Current open-source datasets for user behavior simulation fall short in several key aspects. First, most datasets record only sparse, decontextualized–or even synthetic–user actions. Some shopping datasets like Amazon-M2 or ECInstruct Jin et al. (2024a, b) record only the isolated actions (e.g., purchases or clicks) with limited observation context. Others Deng et al. (2024); Chen et al. (2024); Yao et al. (2022) use synthetic or third-party annotated behaviorial data, which lacks the individual behavior pattern and the authenticity. Additionally, few datasets provide step-level reasoning or persona information, despite prior work has shown that rationale can improve LLM agent’s performance in behavior and decision modeling Lu et al. (2025a); DeepSeek-AI et al. (2025). Similarly, user persona strongly correlates with behavioral patterns Helmi et al. (2023), making persona essential for durable personalization experience.
To address these limitations, we introduce OPeRA, a dataset of Observation, Persona, Rationale, and Action collected from real human users during online shopping. OPeRA provides rich, time-aligned logs of users’ web browsing behavior, completed with self-reported rationales and detailed self-reported persona profiles. Unlike prior datasets, OPeRA captures not only what users do but also why they do it, enabling deeper insights into decision-making processes. This paper focus on the online shopping domain as a beginning point due to its everyday prevalence, complex decision flow, and strong ties to personalization Mican and Sitar-Taut (2020); Wei (2016). Online e-commerce environments like Amazon requires a user’s multi-step interactions involving comparisons, trade-offs, and goal-directed behaviors in one shopping session, all of which are prime testbeds for studying the capacity of LLMs to simulate real human actions.
To collect the OPeRA dataset, we developed ShoppingFlow, a custom browser plugin that captures user interactions alongside corresponding web context and triggers rationale prompts at decision points, as shown in Figure 1. We also collect rich persona information through an online survey and an optional interview to include user profile information such as demographics, shopping styles, and personality traits.
The OPeRA-full contains 692 shopping sessions from 51 unique users, 28,904 <action, observation> pairs, and 604 human-annotated rationales. After post-processing, we also provide OPeRA-filtered, which includes 527 sessions, 5856 <observation, action> pairs and 207 rationales. OPeRA serves as the first benchmark dataset for evaluating LLM agents on personalized and verifiable user behavior simulation. We benchmark four state-of-the-art LLMs (GPT-4.1 OpenAI (2025a), DeepSeek-R1 DeepSeek-AI et al. (2025), Claude-3.7 Anthropic (2025), and Llama-3.3 Meta (2024)) on OPeRA-test (a subset of the OPeRA-filtered) and analyze their ability to predict the next action and rationale of a specific user based on their persona and interaction history. These findings lay a foundation for future work on building LLM-powered digital twins capable of accurate and adaptive behavior modeling.
{booktabs}
[…]
ECInstruct-SA 10k Sentiment Analysis ✗ ✗ ✗ Purchase User
ECInstruct-REC 10k Recommendation ✗ ✗ ✗ Purchase User
Amazon-M2 3M Recommendation ✗ ✗ ✗ Click User
Repeat Buyers 54M Buyer Prediction ✗ ✓ ✗ Click, Cart
Favor, Purchase User
Taobao 100M Recommendation ✗ ✗ ✗ Click, Cart
Favor, Purchase User
YOOCHOOSE 9M Purchase Prediction ✗ ✗ ✗ Click, Purchase User
Shopping MMLU 3973 Recommendation ✗ ✗ ✗ Query, Click
Purchase User
[…]
Type, Select Annotator,
GPT
GUI-WORLD 12k GUI Understanding
Instruction Follow ✓ ✗ ✗ Click, Paste,
Search, Type Annotator,
Video
WebArena 812 Web Navigation
Instuction Follow ✓ ✗ ✗ Click, Hover,
Type, Tab Switching,
Navigation Annotator,
GPT
WebShop 1600 Web Navigation ✓ ✗ ✗ Input, Click Annotator
[…]
OPeRA-filtered 692
527 All Above and
User Behavior Simulation ✓ ✓ ✓ Basic Action,
Semantic Action User
[…]
Table 1: Properties of existing datasets compared to OPeRA. “O”: Environment Observation. “Pe”: User Persona. “R”: Rationale behind action. “A”: Action Space. “Source”: Action Source.
2 Related Works
2.1 LLMs for Human Behavior Simulation
Large Language Model agents can handle complex tasks Wang et al. (2024b); Yao et al. (2023); Shinn et al. (2023); Wu et al. (2023), and researchers are using them as human proxies across domains, from social-science simulations Park et al. (2023); Wang et al. (2024a); Sreedhar et al. (2025) and recommender-system evaluation Wang et al. (2023) to UX testing Lu et al. (2025b) and health-counsellor training Louie et al. (2025). They even reproduce classic results in experimental psychology and economics, such as Milgram Shock Experiment Aher et al. (2023).
In parallel, there have been works looking at generating synthesized personas based on text data, such as PersonaHub Ge et al. (2024) and the approach by  Shi et al. (2025), demonstrating promise in downstream modeling. Moreover, several approaches have integrated persona information to enrich behavioral simulation Shao et al. (2023); Chuang et al. (2024); Shi et al. (2025). For example, Park et al. (2024) introduces an persona-grounded framework using qualitative interviews, enabling agents to accurately simulate individual preferences, attitudes, and behaviors, showing that incorporating personas improves the realism of simulated behavior.
Moreover, there is a growing trend for using LLM agents in online scenarios to produce human-like behaviors, such as Claude Computer Use Anthropic (2024), AutoGLM Liu et al. (2024), Coco-Agent Ma et al. (2024), Mobile-Agent-E Wang et al. (2025b), and OpenAI Operator  OpenAI (2025b), enabling interaction in more complex human-computer settings. Yet, existing research on LLM web agents predominantly focuses on optimizing and evaluating agents for task completion Gur et al. (2023); Zhou et al. (2024a); He et al. (2024), training agents to learn the fewest, most direct steps–whereas real users behaviors normally contain richer and nonlinear paths. There remains a lack of work exploring accurate simulation of human behavior, particularly the modeling of personalized user behaviors. Motivated by this limitation, we propose OPeRA, which aims to better facilitate research on simulating realistic and personalized human behaviors in online scenarios.
Figure 2: Pipeline of our Chrome Plugin, ShoppingFlow. Content Script detects click, scroll and input actions. Background Script detects page-related actions and handles data uploading. Rationale pop-up is triggered at a certain probability when certain action types are detected.
2.2 User Online Behavior Datasets
Existing datasets containing user online behaviors can be broadly categorized into two sets, shown in Table 1. The first category consists of recommendation-oriented datasets which mostly records user interaction with product items (e.g., click, purchase), lacking of essential context. A commonly used example is the Amazon Review 2023 dataset Hou et al. (2024), which contains over 571 million user review-writing behaviors and links them with product information. Follow-up studies, such as ECInstruct (Peng et al., 2024), have utilized the Amazon Review to work on tasks like user review sentiment analysis. Others, such as Amazon-M2 Jin et al. (2024a), YOUCHOOSE Ben-Shimon et al. (2015), Repeat Buyers Liu et al. (2016), and Taobao Zhu et al. (2018), capture richer shopping behaviors such as clicks, add-to-cart, and purchases. The Repeat Buyers dataset further offers basic user persona like gender and age. However, these datasets lack fine-grained user behaviors and contextual information needed to explain how actions and decisions are made at each stage of the user shopping journey, limiting their utility in user behavior simulation.
A separate line of datasets focuses on task completion, with environment observations for web agents tasks, such as Mind2Web Deng et al. (2024), GUI-WORLD Chen et al. (2024), WebArena Zhou et al. (2024b), and WebShop Yao et al. (2022). These datasets contain user interaction traces on website or mobile devices (e.g., clicks, typing) and HTML or screenshots. While useful for studies like instruction following, these datasets are built from annotator-generated or synthetic interaction logs. As a result, they lack both behavioral authenticity and personalization, limiting their realism.
There are other datasets and benchmarks which focus on evaluating LLM agents’ reasoning ability using question-answering format, for example Mialon et al. (2023). While such tasks are valuable for evaluating models, these datasets are not suitable for simulating human behavior.
In summary, existing datasets capture certain aspects of user behaviors but lack the comprehensive data required for simulating real user behavior.
3 OPeRA Dataset
Our work presents a novel dataset OPeRA to advance NLP research in realistic user behavior simulation, particularly within the context of online shopping scenarios.
3.1 OPeRA Data Collection
3.1.1 Participant Recruitment
A total of $84$ participants were recruited through snowball sampling Goodman (1961). During the pre-screening survey, candidate participants were required to self-identify if they were frequent Amazon customers and if they planned to make at least one purchase on Amazon in the next several weeks. These two inclusion criteria helped secure a reliable data collection process. All participants were required to install the ShoppingFlow Chrome plugin (details in Section 3.1.3), shop normally on the Amazon website for a four-week period, and participate in an online survey and an optional interview. Details of recruitment are in Appendix A.
3.1.2 Persona Information Collection
We collected detailed user persona information through a structured online survey and an optional semi-structured interview. All survey questions are designed based on established work (design details in Appendix B) to solicit consumer characteristics that have been shown to correlate with shopping behaviors. The survey consists of three main sections: demographic information, shopping preferences, and personality traits. Demographic information includes age, gender, education, occupation, income, residence, and self-description. Shopping preferences includes online shopping frequency, membership status, shopping habits, seasonality, advertising trust, review engagement, delivery influence, and an adapted eight-item Consumer Styles Inventory (CSI) Nayeem and Marie-IpSooching (2022). Personality traits are measured using the Big-Five Inventory Goldberg (1992) as well as a self-reported MBTI personality Myers et al. (1998).
Following prior work on persona information collection Park et al. (2024), we invited participants to attend a 20-minute optional semi-structured interview about their demographic information, personal background, and online shopping preferences, which aimed to better contextualize the personalized decision-making processes behind users’ online shopping actions. The interview design description and the protocol used are provided in Appendix C.
Figure 3: OPeRA Dataset Overview. The dataset comprises four major components: action traces, web observations, rationales, and user personas. Each shopping session is a sequence of timestamped actions. Each action is paired with a corresponding web observation, which includes: the full HTML of the interacted webpage, a simplified HTML with key elements, a screenshot, and product information for purchases (if applicable). The rationale is a natural language explanation of why the user performed the action. The persona contains detailed user profiles collected from surveys and interviews, covering demographics, personality traits, and shopping preferences.
3.1.3 Shopping Behavior Collection
To support the construction of the dataset, we designed ShoppingFlow, a Chrome extension that automatically captures user behaviors and contextual web observations during Amazon shopping sessions (shown in Figure 2).
The plugin includes two main scripts: a Content Script that runs within the Amazon page to log user interactions (including inputs, clicks, and scrolls) with timestamps, target elements, and HTML using custom parsing rules (Appendix D); and a Background Script that tracks page-level events like reloads and navigation. All data was securely uploaded to Amazon S3. To capture rationale, the plugin randomly triggers pop-ups (8% chance) asking users to explain their actions (question design in Appendix E).
3.1.4 Post-Processing
To ensure data quality and to protect user privacy, we applied a multi-step post-processing procedure. We configured the plugin to not record any personally identifiable information (PII), such as the user login page, account profile page, or the checkout details. In addition, we designed a rule-based automated detection and pattern matching script to mask any PII unavoidably contained in a page (e.g., username in navigation bar), including usernames, zip codes, addresses, specific workplaces, and payment details, before any human touch. Lastly, we manually checked the data to ensure there is no PII in the dataset.
The actual purchase information is not collected since it is in the checkout detail page. Instead, we infer a “purchase” action via click actions on “proceed to checkout,” “buy now,” and “set subscription” buttons. We associate the inferred purchase action with the corresponding product information during post-processing.
In addition, the raw user data is a stream of continuous user action sequences that do not separate different shopping sessions. Thus, our team segmented user actions using a two-step strategy based on temporal intervals and purchase signals. Sessions were first split using a time threshold and then further segmented at purchase intention events (i.e., clicking on “proceed to checkout / buy now / set subscription / add to cart” button). Detailed explanation about the selection of the threshold can be found in Appendix F. Finally, sessions with fewer than five actions were discarded to remove trivial or non-informative behaviors as a prior work reported meaningful sessions typically contain at least six to seven interactions Wang et al. (2025a).
To reduce noise and improve the quality of behavioral data, some actions that occurred on uncommon or rarely visited pages or do not reflect meaningful intent are removed. Moreover, we filtered out clicks on non-interactive areas such as the background, and further filtered actions involving Amazon Rufus.
To support behavior modeling and evaluation under a more tractable setting, we follow prior work Lu et al. (2025a) and similarly define a simplified action $\mathcal{A}$ space consisting of key interaction types. Specifically, we retain three high-level actions that are both semantically meaningful and commonly observed: input, click, and terminate. Within the “click” category, we further differentiate between several subtypes to reflect distinct user intents, shown in Table  3.2. This abstraction reduces complexity while preserving the core structure of user behaviors, enabling more stable behavior simulation.
3.2 Dataset Details
An overview of the dataset is presented in Figure 3. Of all the users, 51 contributed at least one shopping session. In total, the OPeRA-full contains 692 sessions, 28,904 <action, observation> pairs, and 604 rationale annotations. The OPeRA-filtered contains 527 sessions, 5,856 <action, observation> pairs, and 207 rationales. Table 2 presents the statistics. Table 3.2 shows the click type distribution.
User Persona
For participant $i$, the persona is represented by $P_{i}$, comprising two components: a structured survey and interview. Both focusing on their demographics, personality or personal background, and shopping preferences.
User Action Traces
Each shopping session $j$ includes users’ web interactions on the shopping website. The action trace is represented by $A_{j}=\{a_{1},...,a_{T}\}$, where $a_{t}\in\mathcal{A}$ represents the user’s action at step $t$ in the action space $\mathcal{A}$. The Action space $\mathcal{A}$ includes click, scroll, input, navigate, tab activate. Each action is assigned a unique identifier (UUID) and is timestamped to preserve the exact temporal order.
In particular, for each click action, the corresponding CSS selector of the element is provided to uniquely identify the clicked target. Additionally, semantic identifiers are assigned to click actions to indicate their functional context, such as interactions with products in search results (e.g., semantic_id: “search_result.product_name”) or interactions with other commonly used page elements. This semantic id enables downstream models to have a clearer recognition and observation of user behaviors when utilizing the dataset. Similarly, scrolling actions are presented with their start and end positions, supporting analyses of how users explore page content.
Rationale
Alongside the user actions, rationales are provided for some actions, $R_{j}=\{r_{1},\dots,r_{T}$, where $r_{k}$ is a nullable string describes the user’s rationale for specific action $a_{k}$, explicitly capturing the underlying motivations or thought processes of users. These insights enable the downstream models to have a thorough understanding of why users make particular choices, offering a deeper view of user behavior and the reasoning process.
Web Observation
In addition to action traces and rationales, the observation of the web context at each action step in session $j$ is captured, represented by $O_{j}=\{o_{0},...,o_{T}\}$. Each $o_{t}$ includes the HTML content and a screenshot [^2] The experiments presented in this paper did not involve the use of screenshot data.. The HTML content also contains annotations indicating the viewability of each clickable element, as well as relevant page metadata (such as product name, product price, etc.) at time $t$. To reduce noise and storage while keeping the HTML structure, we provides the a simplified version of the HTML content containing key elements for frequently encountered pages, such as search results, product detail pages, and shopping carts.
Additionally, if a purchase is made during a session, the final purchase information, including the price, product title, product options, and Amazon Standard Identification Number (ASIN), are recorded to facilitate potential downstream tasks such as recommendation.
{booktabs}
[…]
Click 5,253 (18.1%) 5,051 (86.3%)
Tab Activate 1,945 (6.7%) –
Navigate 1,901 (6.6%) –
Text Input 606 (2.1%) 597 (10.2%)
Terminate – 208(3.6%)
[…]
Table 2: Action type distribution.
{booktabs}
[…]
review 1052 20.8%
search 763 15.1%
product_option 700 13.9%
product_link 537 10.6%
other 449 8.9%
purchase 321 6.4%
nav_bar 283 5.6%
page_related 198 3.9%
quantity 191 3.8%
suggested_term 182 3.6%
cart_side_bar 145 2.9%
cart_page_select 139 2.8%
filter 91 1.8%
[…]
Table 3: Click type distribution in OPeRA-filtered. Detailed descriptions in Appendix G
{booktabs}
[…]
Avg. # of Action 11.11
Avg. # of Input 1.13
Avg. # of Click 9.58
Avg. # of Terminate 0.39
[…]
Table 4: OPeRA-filtered dataset statistics per-session
4 Tasks and Experiments
4.1 Tasks
We show how OPeRA can be leveraged to evaluate LLM’s ability to simulate consumer behavior in online shopping.
Next Action Prediction
The next action prediction task aims to model the user’s next action based on previous behaviors following the definition of previous work (Deng et al., 2024; Lu et al., 2025a). Given a history of actions $\{a_{1},\cdots,a_{t-1}\}$ in shopping session $j$, corresponding web contexts $\{o_{1},...,o_{t}\}$, rationale $\{r_{1},...,r_{t}\}$ [^3] Note: Rationale annotations are sparse, whereas actions and observations are fully recorded at each time step., and the consumer profile $P_{i}$, the model is tasked to predict the immediate next action $a_{t}$, learning a function of the form:
$a_{t}=F_{action}(a_{1...t-1},r_{1...t-1},o_{1...t},P_{i})$
4.2 Experiments
4.2.1 Experiment Setup
From the OPeRA-filtered, we furthur construct the test set OPeRA-test by randomly sampling 15 out of 51 users and randomly sampling 90 sessions from these users, resulting in 992 actions. We evaluate four state-of-the-art LLMs, including two open-source models, Llama-3.3-70B-Instruct and DeepSeek-R1, and two proprietary models, GPT-4.1 and Claude-3.7-Sonnet.
All models are evaluated in a zero-shot, prompt-based setting without fine-tuning. Prompt templates are provided in Appendix I.
To investigate how different input factors affect model behavior, we conducted a series of ablation studies. Specifically, we excluded persona information (w/o persona) and additionally removed the history rationale from the input (w/o rationale).
{booktabs}
[…]
  w/o persona 22.06 45.55 43.45 58.47
   w/o rationale 21.28 34.93 42.63 51.17
[…]
  w/o persona 15.52 27.43 33.86 48.86
   w/o rationale 15.74 27.16 32.65 47.92
[…]
  w/o persona 10.75 25.33 22.76 43.10
   w/o rationale 10.08 26.06 20.29 43.10
[…]
  w/o persona 8.31 23.69 18.59 33.21
   w/o rationale 8.76 23.60 19.22 34.19
[…]
Table 5: Evaluation of actions in next action prediction task. We report four metrics here to assess model performance (Full results can be found in Appendix H): Action Generation Accuracy measures the exact-match accuracy of the predicted next action; Action Type Macro F1 evaluates the model’s ability to predict the correct high-level action category (e.g., input, click, terminate); Click Type Weighted F1 captures the performance of predicting the specific type of click actions; Session Outcome Weighted F1 reflects how well the model can predict the final outcome of a session, where each session ends either in a purchase or a terminate action. “Claude-3.7”: Claude-3.7-Sonnet, “Llama-3.3”: Llama-3.3-70B-Instruct. All metrics are reported as percentages (%). Instance size $n=902$.
4.2.2 Evaluation
To assess the accuracy of generated user actions, we apply an exact match criterion: a prediction is correct only when all required components align with the ground truth. Specifically, for click actions, the clicked target name must match. For input actions, this includes identifying the input field and generating exact input text.
In addition to exact match accuracy, we evaluate the model’s ability to classify action types. We report weighted F1 for high-level action types (click, input, terminate). Given the highly imbalanced nature of user behavior distributions, we also report macro F1 to highlight performance across all classes regardless of frequency.
To further examine fine-grained prediction capabilities, we evaluate the weighted F1 score for click subtypes. This captures whether the model not only predicts that a user will click but also understands the specific type of click (e.g., review, product_link, purchase).
Finally, given the goal-oriented nature of online shopping, we assess the model’s ability to predict session outcomes. Each session in the ground truth ends in either a click on purchase-related button or a terminate action. We evaluate the model’s accuracy and F1 score on these terminal actions to understand whether it can correctly capture users’ long-term goals and decision-making processes.
4.2.3 Results and Analysis
The results for next action prediction are presented in Table 6. Among all models, GPT-4.1 achieves the strongest overall performance. It obtains the highest action generation accuracy of 22.06%, a macro F1 score of 48.78% on action type prediction, and a weighted F1 score of 44.47% on click type prediction. For session outcome prediction, GPT-4.1 also shows solid performance with 58.47% F1. This strong performance may be attributed to its large context window, which facilitates better processing of complex interaction histories. DeepSeek-R1 performs moderately well across tasks. It achieves a maximum action generation accuracy of 15.74% and obtains solid outcome prediction, possibly due to its strong reasoning abilities. However, the performance may be limited by the model’s 128k context length. Claude-3.7 shows modest performance. Its action generation accuracy is around 10.75%, and action type prediction is generally weaker than GPT-4.1 and DeepSeek. Nonetheless, it achieves relatively good outcome prediction, suggesting some robustness in capturing high-level user intent. LLaMA-3.3 underperforms across all metrics, with action generation accuracy of only 8.76%. Its lower performance may be due to the smaller model size (70B) and shorter context length (128k).
The role of persona information varies across models. While adding persona information does not consistently improve exact action generation accuracy, it generally enhances the model’s performance on action type and click type prediction across all model families. This suggests that persona information provides useful priors about user preferences and behavior patterns, helping the model better classify action semantics. However, its limited effect on action generation accuracy implies that simply including persona in the prompt may introduce noise rather than help. Current models have limited ability to deeply integrate persona into step-level decision-making. This highlights potential room for improvement in personalized user modeling.
Additionally, removing historical rationales consistently leads to performance degradation across most models and metrics, particularly in outcome prediction. This confirms that rationale information serves as valuable intermediate supervision, guiding the model to align its decisions with plausible user intent. We also note several outliers in the results. For instance, LLaMA-3.3 does not consistently benefit from rationale inputs. This may be due to its smaller model size and limited capacity to leverage additional contextual signals effectively.
5 Conclusion
This paper introduces OPeRA, a comprehensive online shopping behavior dataset specifically designed to advance the development and evaluation of LLM-based agents for simulating user behavior. By capturing full shopping trajectories, including action traces, web observations, user personas, and explicit rationales, the dataset provides a verifiable, personalized resource for user behavior modeling. We define a suite of evaluation tasks and conduct a comprehensive analysis across four state-of-the-art LLMs. Our results highlight both the promise and current limitations of LLM agents in simulating realistic user behavior. While certain models demonstrate plausible rationale and action prediction under simple setups, there remains substantial room for improvement, especially in handling complex decision flows and deeper personalization.
Limitations
This study evaluates LLM agents under a simplified setup, following prior work Lu et al. (2025a), adopting a reduced action space and a coarse-grained session segmentation strategy. In particular, actions such as scrolling and page navigation are omitted. This simplification serves to manage the complexity of the task: while text-based LLMs are robust in processing structured language data, they might struggle to accurately model continuous UI-based actions like scrolling. As a result, the simplified action traces become clearer and more suitable for preliminary analysis. We hope that future models capable of richer, multimodal reasoning may better handle these complexities.
Second, although we collect screenshots alongside HTML data for every user interaction, the experiments in this paper do not incorporate visual signals. This is primarily because existing LLM agents are not yet robustly equipped to interpret raw visual UI elements from screenshots in conjunction with structured web content. We envision that future multimodal agents can leverage this visual information to better capture UI layouts, visual salience, and non-textual cues, which are often crucial for interpreting user decisions.
Looking forward, the OPeRA dataset enables a wide range of future research directions beyond those explored in this paper. For example, the dataset can support downstream tasks such as purchase prediction and user intent inference, benefiting recommender systems by grounding them in real, rationalized user behavior. Furthermore, simulation based on OPeRA offers a scalable framework for generating synthetic interaction data, which can be valuable for applications like website design and adaptive user interaction. We hope this dataset serves as a foundation for more realistic, personalized, and interpretable user behavior modeling in the LLM era.
Appendix A Data Collection Implementation Details
We recruited 84 participants via snowball sampling. The participants are pre-screened to ensure that they meet the following criteria: at least 18+ years old, based in the U.S., English speaker and have used (or plan to use) Amazon website to make a purchase in the past (future) couple of weeks.
The participant incentive structure consists with: a) \$5 for completing the online survey (10 mins) and b) \$10 for participating in an optional 20-mins interview to discuss about their personal background. c) \$5 per one qualified week for participants who have complete one or more purchase sessions, with at least two rationale recorded by the ShoppingFlow plugin. In addition, we provide \$10 as a bonus to participants who can successfully complete four or more qualified weeks of data collection in a row. All incentives were delivered as Amazon digital gift cards to their emails.
During the data collection process, no personally identifiable information was retained. To ensure data privacy, we don’t collect data on sensitive pages like checkout or account page. Meanwhile, we implemented a script to anonymize sensitive details in recorded context data, such as zip codes, names, and addresses. Any screenshots that unintentionally contained identifiable information were reviewed and removed prior to dataset release. Furthermore, a research assistant conducted a post-screening review to confirm the exclusion of identifiable information.
This study was conducted in compliance with the Institutional Review Board (IRB) guidelines at Northeastern University.
Appendix B Survey Design
The survey consists of three main sections: demographic information, shopping preferences, and personality traits.
Demographic information significantly influences consumer behavior Hou and Elliott (2021). The survey collects age, gender, education level, occupation, family income, location of residence, and a two-sentence self-description.
Shopping preferences section asks for participants’ online shopping frequency and whether they have a paid membership or not. In addition, we include 12 questions inspired by previous literature, all with a 5-point Likert scale answer Likert (1932). These questions include four shopping habits items, seasonality SpendingPulse (2023), tendency of believing in advertisements Nizam and Jaafar (2018), habits of reading product reviews Shao et al. (2014), and the influence of delivery Baubonienė and Gulevičiūtė (2015) and an 8-items consumer styles inventory (CSI) adapted to the online shopping context  Sprotles and Kendall (1986); Helmi et al. (2023); Prakash et al. (2018); Nayeem and Marie-IpSooching (2022), including: brand loyalty, price conciousness, perfectionism and high-quality conciousness, impulsiveness, confusion by overchoice, brand consciousness, recreational consciousness, and novelty and fashion conciousness.
Personality traits section utilizes the Big-Five Personality Inventory with five core dimensions: Openness to experience, conscience, extrovertism, agreeableness, and neuroticism Goldberg (1992). A self-reported MBTI is also included Myers et al. (1998).
B.1 Survey Questions
B.1.1 Demographic Information
Q1: Gender.
[Male; Female; non-binary]
Q2: Age.
[Under 18; 18-24; 25-34; 35-44; 45-54; 55-64; 65+]
Q3: Which city do you live?
Q4: What is your highest level of education ?
[High school diploma or lower; Bachelors’ degree or current college student; Graduate degree or current grad student (MA, MS, MBA, etc.); Doctoral degree or current doctoral student (PhD, JD, MD, DDS etc.); Prefer not to say]
Q5: Do you live alone or live together with others, if so who are they? (Optional)
Q6: What was your household before-tax income during the past 12 months? If you are a student, what’s your allowance or stipend?
[Less than \$25,000; \$25,000-\$49,999; \$50,000-\$74,999; \$75,000-\$99,999; \$100,000-\$149,999; \$150,000 or more; Prefer not to say]
Q7: What best describes your employment status over the last three months?
Q8: What best describes your employment status over the last three months?
[Full-time employee; Part-time employee; Self-employed; Unemployed and looking for work; Student; Retired; Other:
Q9: Use two sentences to describe yourself. Example 1:“I am a machine learning researcher at a startup focusing on autonomous driving. My daily work include developing and optimizing deep learning models for perception, sensor fusion, and decision-making in self-driving vehicles.” Example 2: “I am a PhD student in computer science, specializing in AI for healthcare. I frequently conduct experiments, analyze medical datasets, and collaborate with doctors to ensure our models are clinically interpretable. I also attend academic seminars, present my findings at conferences, and participate in lab meetings to refine research directions. ”
B.1.2 Shopping Preferences
Q10: How often do you shop online?
[More than three times a week; Once to twice a week; Once every couple of weeks; Less than once a month]
Q11: How much money (in US dollars) do you spend on online shopping per month? (Not including food or delivery services)
Q12: Do you have a paid Amazon Prime membership?
[Yes; No]
From Q13 to Q24, all items use the same response scale: [Strongly disagree, Somewhat disagree, Neither agree nor disagree, Somewhat agree, Strongly agree]
Q13: I tend to shop more during holidays (e.g. Black Friday, holiday sales).
Q14: Online ads attract my attention and are a good source of information.
Q15: I usually do a lot of research (e.g. reading online reviews) before making a purchase.
Q16: I prioritize delivery speed and delivery fee of the product.
Q17: Getting high-quality online products is very important for me.
Q18: The more expensive online product brands are usually my choice.
Q19: The more I learn about online products, the harder it seems to choose the best.
Q20: I shop quickly for online products, buying the first product or brand I find that seems good enough.
Q21: Once I find a brand I like, I stick with it.
Q22: I would buy a new or different brand of product just to see what it is like.
Q23: I enjoy shopping for online products just for the fun of it.
Q24: I look carefully to find the best value for money when shopping online.
B.1.3 Personality Traits
Big Five Test: read the statements carefully and indicate to what extent you agree of disagree. From Q25 to Q74, all items use the same response scale: [Very Inaccurate, Moderately Inaccurate, Neither Accurate Nor Inaccurate, Moderately Accurate, Very Accurate]
Q25: Am the life of the party.
Q26: Feel little concern for others.
Q27: Am always prepared.
Q28: Get stressed out easily.
Q29: Have a rich vocabulary.
Q30: Don’t talk a lot.
Q31: Am interested in people.
Q32: Leave my belongings around.
Q33: Am relaxed most of the time.
Q34: Have difficulty understanding abstract ideas.
Q35: Feel comfortable around people.
Q36: Insult people.
Q37: Pay attention to details.
Q38: Worry about things.
Q39: Have a vivid imagination.
Q40: Keep in the background.
Q41: Sympathize with others’ feelings.
Q42: Make a mess of things.
Q43: Seldom feel blue.
Q44: Am not interested in abstract ideas.
Q45: Start conversations.
Q46: Am not interested in other people’s problems.
Q47: Get chores done right away.
Q48: Am easily disturbed.
Q49: Have excellent ideas.
Q50: Have little to say.
Q51: Have a soft heart.
Q52: Often forget to put things back in their proper place.
Q53: Get upset easily.
Q54: Do not have a good imagination.
Q55: Talk to a lot of different people at parties.
Q56: Am not really interested in others.
Q57: Like order.
Q58: Change my mood a lot.
Q59: Am quick to understand things.
Q60: Don’t like to draw attention to myself.
Q61: Take time out for others.
Q62: Shirk my duties.
Q63: Have frequent mood swings.
Q64: Use difficult words.
Q65: Don’t mind being the center of attention.
Q66: Feel others’ emotions.
Q67: Follow a schedule.
Q68: Get irritated easily.
Q69: Spend time reflecting on things.
Q70: Am quiet around strangers.
Q71: Make people feel at ease.
Q72: Am exacting in my work.
Q73: Often feel blue.
Q74: Am full of ideas.
Q75: What is your MBTI personality type? (Optional)
Appendix C Interview Design
The interview includes question sections of demographic info, detailed personal background, and online shopping habits and preferences. The interviews encourages participants to elaborate on their personas through open responses, thus facilitating a more nuanced understanding of their individual experiences and decision-making processes. For example, in the personal narrative section, participants were encouraged to describe a typical day and share how they perceive themselves. Similarly, in the online shopping-related section, they were asked to describe a recent purchase session on Amazon, providing concrete examples of their shopping habits for downstream models’ understanding and simulation of consumer behavior. This includes pre-purchase research activities, in-session shopping behaviors, engagement with review content, and attitude on advertisement.
C.1 Interview Protocal
Demographics
Introduction:
Can you tell me a bit about yourself? What kind of work do you do? Where do you live? Do you live alone or with family?
Personal Background
Daily Life:
You mentioned your work/study. What does a typical day look like for you?
Work Activities: Can you tell me more about [job/study]? What are your main responsibilities or daily tasks?
After-Work Activities: What do you usually do after work?
Self-Perception: How would you describe yourself?
Online Shopping Preferences
Recent Purchase:
Can you recall a recent Amazon purchase? What was the reason for shopping? What were you looking for? How did you find and decide on the product?
Pre-Shopping Activities: Did you research before making the purchase?
During-shopping Shopping: How long did it take to decide? How many products did you compare before choosing?
Decision Factors: What mattered most in your choice? Do your priorities change based on category (e.g., style for clothing, brand for electronics, price for essentials)?
Reviews: Did you read any reviews before purchasing? What information were you looking for? If not, why?
Advertisements: Do you notice sponsored products? How do ads influence your decisions?
Appendix D Web Parser Design
This section introduces the parser designed to process and simplify web pages, enabling downstream LLMs to better understand and interact with the page content. The parser is guided by a framework called a recipe, which consists of a set of JSON-based rules tailored to specific web pages. These recipes are created through a combination of manual rule-writing and automated assistance from GPT.
Each recipe uses CSS selectors to identify and extract key HTML elements that are important for user interaction (e.g., search boxes, filter options) or containing important semantic information (e.g., product prices, availability status).
We designed recipes for several common page types, such as search result pages, product detail pages, and checkout pages. When parsing a page, the parser extracts these key elements and annotates them with a unique semantic ID. For instance, the ID refinements.colors.red denotes an element that filters results to red-colored items when clicked. These semantic IDs help the model understand both the structure and the function of the elements. Clickable elements are further annotated with visual markers to inform the downstream model of their interactivity.
To simplify the HTML, the parser removes all irrelevant content and styling, retaining only the extracted key elements. This results in a clean, minimal HTML structure that is easier for the model to analyze and reason over.
Appendix E Rationale Prompt Design
The plugin employs context-aware pop-up questions to collect user rationale across different interaction types. The question design follows a hierarchical structure based on event types and specific interaction targets.
Click Events
  • Click on subscription setup button: You clicked on the set up now button. Can you tell us why you subscribed to this product?
  • Click on buy now button: You clicked on the "buy now" button. Why did you do that?
  • Click on add to cart button: You clicked on the "add to cart" button. Why did you decide to add this product to your cart?
  • Click on search button: You clicked on the "search" button. Why did you make this search?
  • Click on filters: You clicked on this filter. Why did you use this filter?
  • Click on product options: You clicked on this product option. Why did you click this product option?
  • Click on checkout button: You clicked on the "checkout" button. What made you decide to checkout?
  • Click on decrease quantity button: You clicked on the “decrease quantity” button. Why did you click this button?
  • Click on increase quantity button: You clicked on the “increase quantity” button. Why did you click this button?
  • Click on product list: You clicked on this product. Why did you click on this product?
  • Click on other area: We noticed that you just had a click action. Why did you do that?
Scroll Events
  • Page scrolling: We saw that you scrolled up/down this page. What are you looking for?
Navigation Events
  • Back / forward navigation: Why did you decide to return to this page?
Tab Switch Events
  • Tab activation: Why did you leave and come back to this tab?
Appendix F Segmentation Interval Threshold Justification
To determine the time interval threshold for session segmentation, we analyzed the distribution of inter-action time gaps from the first week of user activity. From observation, approximately 98% of intervals were shorter than 4 minutes, and 99% were shorter than 78 minutes. Choosing the 99th percentile (78 minutes) as the threshold balances session granularity, avoiding both excessive fragmentation and excessively long sessions that mix unrelated behaviors.
In addition, behavioral analysis showed that users typically interact with the site for 50 to 150 actions before reaching a natural session boundary, such as a purchase or exit, which means around every 50 to 150 interactions, there is probably a session termination or purchase signal. This observation supports the chosen 99% threshold as both statistically and behaviorally grounded.
Appendix G Click Type Description
This section provides detailed descriptions for the click type categories used in the OPeRA dataset, as presented in Table 3.2. The following categories were established through empirical analysis of user click action distributions.
review (20.8%): Clicks on review-related elements, including review images, star ratings, review filters, etc.
search (15.1%): Clicks on search button or search box.
product_option (13.9%): Clicks on product options such as size selectors, color options etc.
product_link (10.6%): Clicks on product images, product titles, or product links that navigate users to product detail pages.
other (8.9%): Miscellaneous clicks that do not fall into the above predefined categories.
purchase (6.4%): Clicks on purchase-intention elements including “Add to Cart”, “Buy Now”, “Subscribe”, and “Checkout”.
nav_bar (5.6%): Clicks on navigation bar elements such as category menus, amazon logo etc.
page_related (3.9%): Clicks on pagination controls, carousel navigation buttons that control page content display.
quantity (3.8%): Clicks on quantity adjustment controls including increase/decrease buttons and item deletion buttons.
suggested_term (3.6%): Clicks on search suggestions.
cart_side_bar (2.9%): Clicks on elements in shopping cart sidebar.
cart_page_select (2.8%): Clicks on selection elements such as item checkbox in the cart page.
filter (1.8%): Clicks on filtering elements including price filters, brand filters, rating filters, and other product refinement controls.
Appendix H Next Action Prediction Experiment Results
Table 6 shows the full results of next action prediction task.
{booktabs}
[…]
  w/o persona 22.06 82.32 45.55 43.45 55.55 58.47
   w/o rationale 21.28 83.13 34.93 42.63 53.33 51.17
[…]
  w/o persona 15.52 81.72 27.43 33.86 56.67 48.86
   w/o rationale 15.74 81.66 27.16 32.65 53.33 47.92
[…]
  w/o persona 10.75 82.28 25.33 22.76 50.00 43.10
   w/o rationale 10.08 81.08 26.06 20.29 47.78 43.10
[…]
  w/o persona 8.31 78.59 23.69 18.59 28.89 33.21
   w/o rationale 8.76 80.23 23.60 19.22 31.11 34.19
[…]
Table 6: Evaluation of actions in next action prediction task. “Claude-3.7”: Claude-3.7-Sonnet, “Llama-3.3”: Llama-3.3-70B-Instruct. All metrics are reported as percentages (%). Instance size $n=902$.
Appendix I Experiment Prompt Design
Below are the two prompts for action prediction task and joint rationale and action generation task:
PROMPT_FOR_ACTION_PREDICTION = """
<IMPORTANT>
Your task is to predict the immediate next action of a shopper.
You need to pretend that you are a real user shopping on amazon.com.
The history action, rationale, context and the user persona will be provided to you.
Ensure your prediction follows natural behavior sequences (e.g., users may click a search box before typing, type a query before clicking search)
</IMPORTANT>
# Action Space
An action is represented in JSON format, and there are four primary types of actions:
#### 1. input‘:
Type text into an input field.
"type": "input",
"name": "input_name",
"text": "input_text"
#### 2. click‘:
Click on a button or clickable element identified by name‘.
"type": "click",
"name": "clickable_name",
#### 3. terminate‘:
When you are unsatisfied with the current search result and you dont want to buy anything, use terminate to indicate that you want to close the browser window and terminate the task.
"type": "terminate"
# Rationale
Rationale is the reason why the user takes the action. Some of the rationale is provided to you.
# Context
Your context will be the HTML of the amazon page you are looking at. Some interactable elements will be added a unique "name" attribute, which you can use to identify the element to interact with (click or input).
# Persona
The user persona reflects the users demographics, personality, and shopping preference. First identify which aspects of the persona might be relevant to the current shopping context, then consider them only if they naturally align with the ongoing shopping journey. DO NOT RELY ON IT.
# Output Format
You need to predict the next action. Your output should follow a strict JSON format:
"type": "<type>",
<IMPORTANT>
OUTPUT A SINGLE JSON OBJECT, NOTHING ELSE.
</IMPORTANT>